# Startup Acquisitions: Acquihires and Talent Hoarding

### v5

@@ 1 Introduction
Our second main result concerns the effect of acquihires on consumer surplus. While increasing the efficiency of the acquirer, an acquihire leads to a loss of surplus induced by the disappearance of the startup. Assuming that part of the efficiency gain is passed on to consumers, the overall effect of the acquihire on consumer surplus is ambiguous. In particular, if high-match acquisitions increase it while low-match acquisitions decrease it, competition authorities who cannot identify match quality may face complex challenges in regulating acquihires. For instance, the welfare effects of banning acquihires might not depend straightforwardly on the ex-ante likelihood of a high-match deal. As we show, prohibiting acquihires tends to decrease consumer welfare when that probability is either very high or very low. In the former case, even though low-match firms have a strong incentive to hoard talent, the potential negative impact rarely materializes. In the latter case, low-match firms endogenously choose not to engage in (expensive) talent hoarding, so all observed acquihires are with high-match firms and thus welfare enhancing. It is when the probability of a high match is intermediate – such that low-match firms remain tempted to hoard talent and are not rare – that banning acquihires has the largest scope for enhancing expected consumer welfare.
An important question is the extent to which the emergence of talent hoarding and its welfare implications hinge on our modeling assumptions. To this end, Section 4 considers several variations of our model including changes to the order and timing of moves, bargaining power, or the market’s competitiveness. The key takeaway is that talent hoarding is not an artifact of our model but that the preemption motives and their adverse effects are a robust phenomenon. However, the section’s results also highlight that there are situations in which inefficient acquihires are less of a concern and do not necessarily warrant regulatory attention.
@@
We close the introduction by situating acquihires within the broader context of labor market dynamics and hiring practices. Talent hoarding, as described in our model, is most likely to arise when firms can identify high-impact teams before their competitors. This mechanism could, in principle, extend beyond acquihires to other forms of hiring, such as direct recruitment of employees with scarce and desirable skills. However, we focus on acquihires but not hiring in general because they present a particularly acute setting for talent hoarding: startups provide a clear and verifiable signal of entrepreneurial ability, while the individuals behind them may not yet be widely recognized within the industry. That said, firms may also choose to bypass the acquisition process altogether by directly poaching a startup’s employees (see Bar-Isaac et al., 2023, and the pertaining discussion in the literature review). Indeed, this practice has recently been documented in the context of AI startups and been portrayed as a way to avoid regulator scrutiny that would ensue with an acquisition of the startup (Deutscher, 2024; Heath, 2024). For example, in the case of Microsoft poaching employees from the AI startup Inflection, the Competition and Markets Authority in the UK investigated the transaction and treated it as a merger. Such hiring events can have similar economic and competitive implications to acquihires. Indeed, as the Microsoft/Inflection case suggests, these events may affect the acquirer, the hired employees, and the startup in ways comparable to an acquihire, warranting equivalent regulatory treatment (Madhumita and Hodgson, 2024).
@@
Our paper is most closely related to the literature studying the economics of startup acquisitions. Much of the early literature examined, in various settings, how the prospect of an acquisition affects the incentives of startups and incumbents to invest in innovation (e.g., Mason and Weeds, 2013; Norbäck and Persson, 2012; Gans and Stern, 2000; Phillips and Zhdanov, 2013; Rasmusen, 1988). Following Cunningham et al. (2021), who demonstrated that incumbents may acquire startups for anticompetitive reasons, a large literature has studied the effects a more restrictive merger policy would have on innovation and overall welfare (Motta and Peitz, 2021; Cabral, 2020, 2023; Katz, 2021; Letina et al., 2024). Others examine how acquisitions can steer the direction of innovation (Bryan and Hovenkamp, 2020a; Callander and Matouschek, 2022; Dijk et al., 2024), and Fumagalli et al. (2023) consider the impact of financial constraints. Several papers consider dynamic incentives (Cabral, 2018; Bryan and Hovenkamp, 2020a; Hollenbeck, 2020; Denicolo and Polo, 2021). A key insight is that if the incumbent pulls too far ahead in the technology space, the pace of innovation will go down. There is also the possibility that the incumbent creates a kill-zone that disincentivizes entry, either by acquiring entrants, copying their products, or investing heavily in innovation (Kamepalli et al., 2021; Teh et al., 2022; Shelegia and Motta, 2021; Bao and Eeckhout, 2023). On the empirical side, Ederer and Pellegrino (2023) show that startups increasingly favor acquisitions over IPOs as exit strategies. Finally, several papers empirically study acquisitions in the tech sector (Affeldt and Kesler, 2021a, b; Prado and Bauer, 2022; Jin et al., 2023; Gugler et al., 2023; Eisfeld, 2022; Gautier and Lamesch, 2021).
@@
Our paper also relates to the empirical literature that directly studies acquihires. Ouimet and Zarutskie (2020), Ng and Stuart (2021), Chen et al. (2021), and Chen et al. (2023) show that acquiring talent is indeed an important motivation for acquisitions. However, acquihired employees separate at a higher rate than regularly hired employees (Ng and Stuart, 2021; Verginer et al., 2022), possibly due to a preference for working at startups or a misalignment with the acquirer’s plans (Kim, 2023; Loh et al., 2019). This empirical finding is consistent with our theoretical result.
@@ 5 Hiring and Layoffs
The first period of this expanded model is identical to the baseline model in Section 2: the firms’ match types are private information, and firm 1 has the opportunity to pursue an acquihire before firm 2. Between the periods, the economy suffers a downturn with probability $\delta\in(0,1)$. If a downturn materializes – an event that is publicly observable – the firms may be hit by adverse shocks (see details below). In period 2, the entrepreneur has the option of creating a new startup, once more leading to an outside option of $\pi_{E}$ for her. [^13] There is empirical evidence that acquihired employees who leave the acquirer are likely to join a new startup (Ng and Stuart, 2021; Kim, 2022, 2023). If the entrepreneur was employed by a firm in period 1, that firm must decide whether to continue the relationship (at the cost of $\pi_{E}$) or lay off the entrepreneur, who might then be hired by the other firm. If there was no acquihire in period 1, it is again firm 1 that moves first in period 2.

### v6

@@ 1 Introduction
Our second main result concerns the effect of acquihires on consumer surplus. While increasing the efficiency of the acquirer, an acquihire leads to a loss of surplus induced by the disappearance of the startup. Assuming that part of the efficiency gain is passed on to consumers, the overall effect of the acquihire on consumer surplus is ambiguous. In particular, if high-match acquisitions increase it while low-match acquisitions decrease it, competition authorities who cannot identify match quality may face complex challenges in regulating acquihires. For instance, the welfare effects of banning acquihires might not depend straightforwardly on the ex-ante likelihood of a high-match deal. As we show, prohibiting acquihires tends to decrease consumer welfare when that probability is either very high or very low. When the probability of a high-match deal is high, low-match firms do have a strong incentive to hoard talent but most of the acquihires are with high-match firms, which on average improves welfare. When the probability of a high-match deal is low, low-match firms endogenously choose not to engage in (expensive) talent hoarding, so all observed acquihires are with high-match firms and thus welfare enhancing. It is when the probability of a high match is intermediate – such that low-match firms remain tempted to hoard talent and are not rare – that banning acquihires has the largest scope for enhancing expected consumer welfare.
An important question is the extent to which the emergence of talent hoarding and its welfare implications hinge on our modeling assumptions. To this end, Section 4 considers several variations of our model including changes to the order and timing of moves, bargaining power, or the market’s competitiveness. The key takeaway is that talent hoarding is not an artifact of our model but that the preemption motives and their adverse effects are a robust phenomenon. However, the section’s results also highlight that there are situations in which inefficient acquihires are less of a concern and do not necessarily warrant regulatory attention. In particular, sequential bidding for the startup is important for the negative welfare effects to emerge. If the bidding is simultaneous, or if the order of bids is uncertain, the prevalence of talent hoarding is reduced. Similarly, if the startup enjoys a strong bargaining position, the adverse welfare effects are less likely to materialize. Market structure also plays a role. Talent hoarding is most likely to occur when there are few competitors and especially if there is a dominant firm facing a challenger. It is less likely to occur in a market with a monopoly firm and a potential entrant, or a market with many competitors. Overall, our results suggest that acquihires may be a concern for regulators but perhaps not warrant the same level of scrutiny as more conventional cases.
@@
We close the introduction by situating acquihires within the broader context of labor market dynamics and hiring practices. Talent hoarding, as described in our model, is most likely to arise when firms can identify high-impact teams before their competitors. This mechanism could, in principle, extend beyond acquihires to other forms of hiring, such as direct recruitment of employees with scarce and desirable skills. However, we focus on acquihires but not hiring in general because they present a particularly acute setting for talent hoarding: startups provide a clear and verifiable signal of entrepreneurial ability, while the individuals behind them may not yet be widely recognized within the industry. That said, firms may also choose to bypass the acquisition process altogether by directly poaching a startup’s employees (see Bar-Isaac et al., 2025, and the pertaining discussion in the literature review). Indeed, this practice has recently been documented in the context of AI startups and been portrayed as a way to avoid regulator scrutiny that would ensue with an acquisition of the startup (Deutscher, 2024; Heath, 2024). For example, in the case of Microsoft poaching employees from the AI startup Inflection, the Competition and Markets Authority in the UK investigated the transaction and treated it as a merger. Such hiring events can have similar economic and competitive implications to acquihires. Indeed, as the Microsoft/Inflection case suggests, these events may affect the acquirer, the hired employees, and the startup in ways comparable to an acquihire, warranting equivalent regulatory treatment (Madhumita and Hodgson, 2024).
@@
Our paper is most closely related to the literature studying the economics of startup acquisitions. Much of the early literature examined, in various settings, how the prospect of an acquisition affects the incentives of startups and incumbents to invest in innovation (e.g., Mason and Weeds, 2013; Norbäck and Persson, 2012; Gans and Stern, 2000; Phillips and Zhdanov, 2013; Rasmusen, 1988). Following Cunningham et al. (2021), who demonstrated that incumbents may acquire startups for anticompetitive reasons, a large literature has studied the effects a more restrictive merger policy would have on innovation and overall welfare (Motta and Peitz, 2021; Cabral, 2021, 2024; Katz, 2021; Letina et al., 2024). Others examine how acquisitions can steer the direction of innovation (Bryan and Hovenkamp, 2020a; Callander and Matouschek, 2022; Dijk et al., 2024), and Fumagalli et al. (2023) consider the impact of financial constraints. Several papers consider dynamic incentives (Cabral, 2018; Bryan and Hovenkamp, 2020a; Hollenbeck, 2020; Denicolo and Polo, 2021). A key insight is that if the incumbent pulls too far ahead in the technology space, the pace of innovation will go down. There is also the possibility that the incumbent creates a kill-zone that disincentivizes entry, either by acquiring entrants, copying their products, or investing heavily in innovation (Kamepalli et al., 2022; Teh et al., 2022; Shelegia and Motta, 2024; Bao and Eeckhout, 2023). On the empirical side, Ederer and Pellegrino (2023) show that startups increasingly favor acquisitions over IPOs as exit strategies. Finally, several papers empirically study acquisitions in the tech sector (Affeldt and Kesler, 2021a, b; Prado and Bauer, 2022; Jin et al., 2023; Gugler et al., 2025; Eisfeld, 2024; Gautier and Lamesch, 2021).
@@
Our paper also relates to the empirical literature that directly studies acquihires. Ouimet and Zarutskie (2020), Ng and Stuart (2021), Chen et al. (2021), and Chen et al. (2023) show that acquiring talent is indeed an important motivation for acquisitions. However, acquihired employees separate at a higher rate than regularly hired employees (Ng and Stuart, 2021; Verginer et al., 2022), possibly due to a preference for working at startups or a misalignment with the acquirer’s plans (Kim, 2024; Loh et al., 2019). This empirical finding is consistent with our theoretical result.
@@ 5 Hiring and Layoffs
The first period of this expanded model is identical to the baseline model in Section 2: the firms’ match types are private information, and firm 1 has the opportunity to pursue an acquihire before firm 2. Between the periods, the economy suffers a downturn with probability $\delta\in(0,1)$. If a downturn materializes – an event that is publicly observable – the firms may be hit by adverse shocks (see details below). In period 2, the entrepreneur has the option of creating a new startup, once more leading to an outside option of $\pi_{E}$ for her. [^13] There is empirical evidence that acquihired employees who leave the acquirer are likely to join a new startup (Ng and Stuart, 2021; Kim, 2022, 2024). If the entrepreneur was employed by a firm in period 1, that firm must decide whether to continue the relationship (at the cost of $\pi_{E}$) or lay off the entrepreneur, who might then be hired by the other firm. If there was no acquihire in period 1, it is again firm 1 that moves first in period 2.